Two helical coils of the same length but with different radii are arranged such that one is placed inside the other and both are co-aligned along the same central axis running along their length. They share a mutual inductance which can vary as a function of time by sliding the inner coil out of the outer coil. Calculate the EMF induced in the outer coil at the instant when


i. A current changes through the inner coil at a rate of 5 A s−1 and the mutual inductance is 0.4 mH.

ii. The current through the inner coil is 2 A and changing at a rate of 5 A s−1 and the mutual inductance is 0.4 mH but changing at a rate of 30 mH s−1.

iii. The current through the inner coil is fixed at 2 A and the mutual inductance is changing at a rate of 30 mH s−1.


1
Expert's answer
2021-05-19T20:45:57-0400

given M=mutual inductance

and "i_2" is current in smaller coil.


we know that,


flux through outer coil "(\phi_1)=Mi_2".............1

where M is mutual inductance.


and induced EMF is given by "E=-\frac{d\phi}{dt}" ..................2

putting the values and solving

a) using1 and 2 we get


"E=-M\frac{di}{dt}"

- 0.4mH "\times 5" A/s

= - 2.0mH A /s


b)here using 1 and 2 we get


"E=-(M\frac{di}{dt}+i_2\frac{dM}{dt})"

= -{(0.4 "\times" 5 )+(2 "\times" 30)}

= - 62 mH A/S


c)using 1 qnd 2 we get


"E=-i_2\frac{dM}{dt}"

=-2×30

=-60 mH A/s


(here -ve sign depict direction opposite to original flow of emf in smaller coil )